# thanks: Corresponding author

### v1

@@ IV Conclusions
We have developed and implemented a perturbative triples correction to the relativistic quadratic unitary coupled cluster singles and doubles method (qUCCSD), denoted as qUCCSD[T], within the X2CAMF Hamiltonian framework. The proposed method retains the Hermitian structure of the unitary ansatz while incorporating dynamic correlation effects from triple excitations in a computationally tractable manner. This formulation allows for efficient and accurate treatment of heavy-element systems where both relativistic and correlation effects play a significant role. The triples correction can be straightforwardly derived from the perturbation truncation of the qUCCSDT energy functional
Benchmarking against Full CI and other UCC-based perturbative triples schemes in the non-relativistic regime demonstrates that qUCCSD[T] consistently yields the lowest deviation from exact results. In the relativistic domain, the method performs exceptionally well across a broad set of properties, including ligand dissociation enthalpies, equilibrium bond lengths, harmonic vibrational frequencies, ionization potentials, and electron affinities. In all cases, qUCCSD[T] matches the accuracy of the standard CCSD(T) method, offering a viable alternative with improved theoretical consistency due to its Hermitian form.
These findings position qUCC method as a promising and general-purpose approach for accurate quantum chemical calculations even on classical computers, especially in systems dominated by relativistic effects. Future work will focus on further scaling improvements, extension to excited states, ionized and electron-attached states using a propagator-like approach, and analytic calculation of properties. Work is in progress towards that direction.
@@ V Supplementary Material
The Supplementary Material contains the qUCCSD, qUCCSD(T), CCSD, and CCSD(T) energies for the diatomic and coinage metal test set.

### v2

@@ IV Conclusions
We have developed and implemented a perturbative triples correction to the relativistic quadratic unitary coupled cluster singles and doubles method (qUCCSD), denoted as qUCCSD[T] . The he X2CAMF Hamiltonian framework has been used to include relativistic effect in the calculations. The proposed method retains the Hermitian structure of the unitary ansatz while incorporating dynamic correlation effects from triple excitations in a computationally tractable manner. This formulation allows for efficient and accurate treatment of heavy-element systems where both relativistic and correlation effects play a significant role. The triples correction can be straightforwardly derived from the perturbation truncation of the qUCCSDT energy functional benchmarking against Full CI and other UCC-based perturbative triples schemes in the non-relativistic regime demonstrates that qUCCSD[T] consistently yields the lowest deviation from exact results among existing triples correction schemes to unitary coupled cluster. In the relativistic domain, the method performs exceptionally well across a broad set of properties, including ligand dissociation enthalpies, equilibrium bond lengths, harmonic vibrational frequencies, ionization potentials, and electron affinities. In all cases, qUCCSD[T] matches the accuracy of the standard CCSD(T) method, offering a viable alternative with added advantage due to its Hermitian nature.
The new triples correction scheme makes qUCC method as a promising and general-purpose approach for accurate quantum chemical calculations even on classical computers, especially in systems dominated by relativistic effects. Future work will focus on further scaling improvements, extension to excited states, ionized and electron-attached states using a propagator-like approach, and analytic calculation of properties. Work is in progress towards that direction.
@@ V Supplementary Material
The Supplementary Material contains the bond dissociation enthalpies for the 18 coinage metal-ligand complexes, as well as energies of the corresponding diatomic and atomic systems used to compute IP and EAs., calculated with FNS-CD-X2CAMF-based CCSD, qUCCSD, CCSD(T), qUCCSD[T] methods.